How do you find the area of a rectangle with these dimensions: 5ft 3in x 6ft?
Answer by jim_thompson5910(35256)

1 ft = 12 inches

5*1 ft = 5*12 inches

5 ft = 60 inches


5 ft 3 in = 5 ft + 3 in = 60 in + 3 in = 63 in

So 5 ft 3 in = 63 in

-----------------------------------------------------

1 ft = 12 inches

6*1 ft = 6*12 inches

6 ft = 72 inches

So 6 ft = 72 inches
-----------------------------------------------------

This means that

(5 ft 3 in)*(6 ft) = (63 in)*(72 in) = (63*72)(in*in) = 4536 square inches


So the area is 4536 square inches